# Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations

### v1

@@ Acknowledgements.
We thank Thibault Garel, Pascal Oesch, Irene Shivaei, Charlotte Simmonds, Andrew Hopkins, Daniel Schaerer, and Rashmi Gottumukkala for useful comments and productive discussions. We gratefully acknowledge support from the CBPsmn (PSMN, Pôle Scientifique de Modélisation Numérique) of the ENS de Lyon for the computing resources. Funded by the European Union (ERC, AGENTS, 101076224). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work made extensive use of several open-source software packages, and we gratefully acknowledge the efforts of their authors: numpy (Harris et al., 2020), astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), ipython (Perez & Granger, 2007), and scikit-learn (Pedregosa et al., 2011).References
- Asada et al. (2024) Asada, Y., Sawicki, M., Abraham, R., et al. 2024, MNRAS, 527, 11372
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Behroozi et al. (2019) Behroozi, P., Wechsler, R. H., Hearin, A. P., & Conroy, C. 2019, MNRAS, 488, 3143
- Bouché et al. (2010) Bouché, N., Dekel, A., Genzel, R., et al. 2010, ApJ, 718, 1001
- Bouwens et al. (2015) Bouwens, R. J., Illingworth, G. D., Oesch, P. A., et al. 2015, ApJ, 803, 34
- Carvajal-Bohorquez et al. (2025) Carvajal-Bohorquez, C., Ciesla, L., Laporte, N., et al. 2025, arXiv e-prints, arXiv:2507.13160
- Chisholm et al. (2019) Chisholm, J., Rigby, J. R., Bayliss, M., et al. 2019, ApJ, 882, 182
- Chruślińska et al. (2024) Chruślińska, M., Pakmor, R., Matthee, J., & Matsuno, T. 2024, A&A, 686, A186
- Clarke et al. (2024) Clarke, L., Shapley, A. E., Sanders, R. L., et al. 2024, ApJ, 977, 133
- Cole et al. (2025) Cole, J. W., Papovich, C., Finkelstein, S. L., et al. 2025, ApJ, 979, 193
- Covelo-Paz et al. (2025) Covelo-Paz, A., Giovinazzo, E., Oesch, P. A., et al. 2025, A&A, 694, A178
- Cullen et al. (2019) Cullen, F., McLure, R. J., Dunlop, J. S., et al. 2019, MNRAS, 487, 2038
- Curti et al. (2024) Curti, M., Maiolino, R., Curtis-Lake, E., et al. 2024, A&A, 684, A75
- Davies et al. (2019) Davies, L. J. M., Lagos, C. d. P., Katsianis, A., et al. 2019, MNRAS, 483, 1881
- De Looze et al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., et al. 2014, A&A, 568, A62
- Endsley et al. (2024) Endsley, R., Stark, D. P., Whitler, L., et al. 2024, MNRAS, 533, 1111
- Ferland et al. (2017) Ferland, G. J., Chatzikos, M., Guzmán, F., et al. 2017, Rev. Mexicana Astron. Astrofis., 53, 385
- Flores Velázquez et al. (2021) Flores Velázquez, J. A., Gurvich, A. B., Faucher-Giguère, C.-A., et al. 2021, MNRAS, 501, 4812
- Fu et al. (2025) Fu, S., Sun, F., Jiang, L., et al. 2025, arXiv e-prints, arXiv:2503.03829
- Götberg et al. (2017) Götberg, Y., de Mink, S. E., & Groh, J. H. 2017, A&A, 608, A11
- Gunawardhana et al. (2013) Gunawardhana, M. L. P., Hopkins, A. M., Bland-Hawthorn, J., et al. 2013, MNRAS, 433, 2764
- Hao et al. (2011) Hao, C.-N., Kennicutt, R. C., Johnson, B. D., et al. 2011, ApJ, 741, 124
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Hayes et al. (2010) Hayes, M., Schaerer, D., & Östlin, G. 2010, A&A, 509, L5
- Hopkins et al. (2001) Hopkins, A. M., Connolly, A. J., Haarsma, D. B., & Cram, L. E. 2001, AJ, 122, 288
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Iyer et al. (2020) Iyer, K. G., Tacchella, S., Genel, S., et al. 2020, MNRAS, 498, 430
- Katz et al. (2023) Katz, H., Rosdahl, J., Kimm, T., et al. 2023, The Open Journal of Astrophysics, 6, 44
- Kennicutt & Evans (2012) Kennicutt, R. C. & Evans, N. J. 2012, ARA&A, 50, 531
- Kennicutt (1998) Kennicutt, Jr., R. C. 1998, ARA&A, 36, 189
- Kroupa (2001a) Kroupa, P. 2001a, MNRAS, 322, 231
- Kroupa (2001b) Kroupa, P. 2001b, MNRAS, 322, 231
- Leitherer et al. (1999) Leitherer, C., Schaerer, D., Goldader, J. D., et al. 1999, ApJS, 123, 3
- Lilly et al. (2013) Lilly, S. J., Carollo, C. M., Pipino, A., Renzini, A., & Peng, Y. 2013, ApJ, 772, 119
- Lin et al. (2025) Lin, X., Egami, E., Sun, F., et al. 2025, arXiv e-prints, arXiv:2504.08028
- Ly et al. (2007) Ly, C., Malkan, M. A., Kashikawa, N., et al. 2007, ApJ, 657, 738
- Mascia et al. (2024) Mascia, S., Pentericci, L., Calabrò, A., et al. 2024, A&A, 685, A3
- Matthee et al. (2022) Matthee, J., Feltre, A., Maseda, M., et al. 2022, A&A, 660, A10
- Matthee & Schaye (2019) Matthee, J. & Schaye, J. 2019, MNRAS, 484, 915
- Meyer et al. (2024) Meyer, R. A., Oesch, P. A., Giovinazzo, E., et al. 2024, MNRAS, 535, 1067
- Michel-Dansac et al. (2020) Michel-Dansac, L., Blaizot, J., Garel, T., et al. 2020, A&A, 635, A154
- Nagaraj et al. (2023) Nagaraj, G., Ciardullo, R., Bowman, W. P., Lawson, A., & Gronwall, C. 2023, The Astrophysical Journal, 943, 5
- Naidu et al. (2024) Naidu, R. P., Matthee, J., Kramarenko, I., et al. 2024, arXiv e-prints, arXiv:2410.01874
- Noeske et al. (2007) Noeske, K. G., Weiner, B. J., Faber, S. M., et al. 2007, ApJ, 660, L43
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Perez & Granger (2007) Perez, F. & Granger, B. E. 2007, Computing in Science and Engineering, 9, 21
- Pirie et al. (2024) Pirie, C. A., Best, P. N., Duncan, K. J., et al. 2024, arXiv e-prints, arXiv:2410.11808
- Rosdahl et al. (2013) Rosdahl, J., Blaizot, J., Aubert, D., Stranex, T., & Teyssier, R. 2013, MNRAS, 436, 2188
- Rosdahl et al. (2022) Rosdahl, J., Blaizot, J., Katz, H., et al. 2022, MNRAS, 515, 2386
- Rosdahl et al. (2018) Rosdahl, J., Katz, H., Blaizot, J., et al. 2018, MNRAS, 479, 994
- Salpeter (1955) Salpeter, E. E. 1955, ApJ, 121, 161
- Scholte et al. (2025) Scholte, D., Cullen, F., Carnall, A. C., et al. 2025, arXiv e-prints, arXiv:2502.10499
- Shim et al. (2009) Shim, H., Colbert, J., Teplitz, H., et al. 2009, ApJ, 696, 785
- Sobral et al. (2013) Sobral, D., Smail, I., Best, P. N., et al. 2013, MNRAS, 428, 1128
- Sparre et al. (2015) Sparre, M., Hayward, C. C., Springel, V., et al. 2015, MNRAS, 447, 3548
- Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
- Steidel et al. (2016) Steidel, C. C., Strom, A. L., Pettini, M., et al. 2016, ApJ, 826, 159
- Sun et al. (2025) Sun, F., Wang, F., Yang, J., et al. 2025, ApJ, 980, 12
- Tacchella et al. (2020) Tacchella, S., Forbes, J. C., & Caplar, N. 2020, MNRAS, 497, 698
- Teyssier (2002) Teyssier, R. 2002, A&A, 385, 337
- Theios et al. (2019) Theios, R. L., Steidel, C. C., Strom, A. L., et al. 2019, ApJ, 871, 128
- Wan et al. (2025) Wan, J. T., Tacchella, S., D’Eugenio, F., Johnson, B. D., & van der Wel, A. 2025, arXiv e-prints, arXiv:2504.05281
- Wang & Lilly (2020) Wang, E. & Lilly, S. J. 2020, ApJ, 892, 87
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Zavala et al. (2021) Zavala, J. A., Casey, C. M., Manning, S. M., et al. 2021, ApJ, 909, 165

### v2

@@ Acknowledgements.
We thank the anonymous referee for the insightful comments that helped improve the manuscript. We also thank Thibault Garel, Pascal Oesch, Irene Shivaei, Charlotte Simmonds, Andrew Hopkins, Daniel Schaerer, and Rashmi Gottumukkala for useful comments and productive discussions. We gratefully acknowledge support from the CBPsmn (PSMN, Pôle Scientifique de Modélisation Numérique) of the ENS de Lyon for the computing resources. Funded by the European Union (ERC, AGENTS, 101076224). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work made extensive use of several open-source software packages, and we gratefully acknowledge the efforts of their authors: numpy (Harris et al., 2020), astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), ipython (Perez and Granger, 2007), and scikit-learn (Pedregosa et al., 2011).References
- Bursty star formation and galaxy-galaxy interactions in low-mass galaxies 1 Gyr after the Big Bang. 527 (4), pp. 11372–11392. External Links: Document, 2310.02314, ADS entry Cited by: §1.
- The Astropy Project: Sustaining and Growing a Community-oriented Open-source Project and the Latest Major Release (v5.0) of the Core Package. 935 (2), pp. 167. External Links: Document, 2206.14220, ADS entry Cited by: §5.
- UNIVERSEMACHINE: The correlation between galaxy growth and dark matter halo assembly from z = 0-10. 488 (3), pp. 3143–3194. External Links: Document, 1806.07893, ADS entry Cited by: §1, §2.
- The JWST UNCOVER Treasury Survey: Ultradeep NIRSpec and NIRCam Observations before the Epoch of Reionization. ApJ 974 (1), pp. 92. External Links: Document, 2212.04026, ADS entry Cited by: §2.
- The Impact of Cold Gas Accretion Above a Mass Floor on Galaxy Scaling Relations. 718 (2), pp. 1001–1018. External Links: Document, 0912.1858, ADS entry Cited by: §1.
- UV Luminosity Functions at Redshifts z 4 to z 10: 10,000 Galaxies from HST Legacy Fields. 803 (1), pp. 34. External Links: Document, 1403.4295, ADS entry Cited by: §1, §4.1.
- Stochastic star formation activity of galaxies within the first billion years probed by JWST. A&A 704, pp. A290. External Links: Document, 2507.13160, ADS entry Cited by: §1.
- Constraining the Metallicities, Ages, Star Formation Histories, and Ionizing Continua of Extragalactic Massive Star Populations. ApJ 882 (2), pp. 182. External Links: Document, 1905.04314, ADS entry Cited by: §4.3.
- Trading oxygen for iron. I. The [O/Fe]-specific star formation rate relation of galaxies. 686, pp. A186. External Links: Document, 2308.00023, ADS entry Cited by: §1.
- The Star-forming Main Sequence in JADES and CEERS at z ¿ 1.4: Investigating the Burstiness of Star Formation. 977 (1), pp. 133. External Links: Document, 2406.05178, ADS entry Cited by: §1.
- CEERS: Increasing Scatter along the Star-forming Main Sequence Indicates Early Galaxies Form in Bursts. 979 (2), pp. 193. External Links: Document, 2312.10152, ADS entry Cited by: §1.
- An H view of galaxy buildup in the first 2 Gyr: Luminosity functions at z 4‑6.5 from NIRCam/grism spectroscopy. 694, pp. A178. External Links: Document, 2409.17241, ADS entry Cited by: §1, §2, §4.1, §4.2.
- The VANDELS survey: the stellar metallicities of star-forming galaxies at 2.5 ¡ z ¡ 5.0. 487 (2), pp. 2038–2060. External Links: Document, 1903.11081, ADS entry Cited by: §3.1.
- JADES: Insights into the low-mass end of the mass-metallicity-SFR relation at 3 ¡ z ¡ 10 from deep JWST/NIRSpec spectroscopy. 684, pp. A75. External Links: Document, 2304.08516, ADS entry Cited by: §1.
- Galaxy And Mass Assembly (GAMA): The sSFR-M∗ relation part I - sSFR-M∗ as a function of sample, SFR indicator, and environment. 483 (2), pp. 1881–1900. External Links: Document, 1811.03712, ADS entry Cited by: §1.
- The applicability of far-infrared fine-structure lines as star formation rate tracers over wide ranges of metallicities and galaxy types. A&A 568, pp. A62. External Links: Document, 1402.4075, ADS entry Cited by: §1.
- The slope and scatter of the star forming main sequence at z~5 : reconciling observations with simulations. arXiv e-prints, pp. arXiv:2510.19044. External Links: Document, 2510.19044, ADS entry Cited by: Figure 4, §4.1.
- The star-forming and ionizing properties of dwarf z 6-9 galaxies in JADES: insights on bursty star formation and ionized bubble growth. 533 (1), pp. 1111–1142. External Links: Document, 2306.05295, ADS entry Cited by: §1.
- The 2017 Release Cloudy. Rev. Mexicana Astron. Astrofis. 53, pp. 385–438. External Links: Document, 1705.10877, ADS entry Cited by: §2.
- REBELS-IFU: dust attenuation curves of 12 massive galaxies at z ≃ 7. MNRAS 539 (1), pp. 109–126. External Links: Document, 2501.10541, ADS entry Cited by: §1.
- The time-scales probed by star formation rate indicators for realistic, bursty star formation histories from the FIRE simulations. 501 (4), pp. 4812–4824. External Links: Document, 2008.08582, ADS entry Cited by: §1.
- Medium-band Astrophysics with the Grism of NIRCam In Frontier Fields (MAGNIF): Spectroscopic Census of H Luminosity Functions and Cosmic Star Formation at z 4.5 and 6.3. 987 (2), pp. 186. External Links: Document, 2503.03829, ADS entry Cited by: §1, §2, §4.1, §4.1.
- Quiescent Low-mass Galaxies Observed by JWST in the Epoch of Reionization. ApJ 954 (1), pp. L11. External Links: Document, 2303.13574, ADS entry Cited by: §4.1.
- Ionizing spectra of stars that lose their envelope through interaction with a binary companion: role of metallicity. 608, pp. A11. External Links: Document, 1701.07439, ADS entry Cited by: §1.
- Galaxy And Mass Assembly: evolution of the H luminosity function and star formation rate density up to z ¡ 0.35. MNRAS 433 (4), pp. 2764–2789. External Links: Document, 1305.5308, ADS entry Cited by: §4.1.
- Dust-corrected Star Formation Rates of Galaxies. II. Combinations of Ultraviolet and Infrared Tracers. 741 (2), pp. 124. External Links: Document, 1108.2837, ADS entry Cited by: §3.1.
- Array programming with NumPy. 585 (7825), pp. 357–362. External Links: Document, Link Cited by: §5.
- The H-alpha luminosity function at redshift 2.2 . A new determination using VLT/HAWK-I. 509, pp. L5. External Links: Document, 0912.3267, ADS entry Cited by: §4.1.
- Toward a Resolution of the Discrepancy between Different Estimators of Star Formation Rate. AJ 122 (1), pp. 288–296. External Links: Document, astro-ph/0103253, ADS entry Cited by: §4.2.
- Matplotlib: a 2d graphics environment. 9 (3), pp. 90–95. External Links: Document Cited by: §5.
- The diversity and variability of star formation histories in models of galaxy evolution. 498 (1), pp. 430–463. External Links: Document, 2007.07916, ADS entry Cited by: §1.
- Impact of metallicity and star formation rate on the time-dependent, galaxy-wide stellar initial mass function. A&A 620, pp. A39. External Links: Document, 1809.04603, ADS entry Cited by: §4.3.
- The SPHINX Public Data Release: Forward Modelling High-Redshift JWST Observations with Cosmological Radiation Hydrodynamics Simulations. 6, pp. 44. External Links: Document, 2309.03269, ADS entry Cited by: §1, §2, §2, §2, §2, §3.1, §3.1, §3.1, §5.
- Star Formation in the Milky Way and Nearby Galaxies. 50, pp. 531–608. External Links: Document, 1204.3552, ADS entry Cited by: §1, §1, §3.
- Star Formation in Galaxies Along the Hubble Sequence. 36, pp. 189–232. External Links: Document, astro-ph/9807187, ADS entry Cited by: §1, §4.1, §4.1, §5.
- On the variation of the initial mass function. 322 (2), pp. 231–246. External Links: Document, astro-ph/0009005, ADS entry Cited by: §2, §3.1.
- Ly Radiative Transfer with Dust: Escape Fractions from Simulated High-Redshift Galaxies. ApJ 704 (2), pp. 1640–1656. External Links: Document, 0907.2698, ADS entry Cited by: §2.
- Starburst99: Synthesis Models for Galaxies with Active Star Formation. 123 (1), pp. 3–40. External Links: Document, astro-ph/9902334, ADS entry Cited by: §1.
- Gas Regulation of Galaxies: The Evolution of the Cosmic Specific Star Formation Rate, the Metallicity-Mass-Star-formation Rate Relation, and the Stellar Content of Halos. 772 (2), pp. 119. External Links: Document, 1303.5059, ADS entry Cited by: §1.
- The Luminosity Function and Clustering of H Emitting Galaxies at z ≍ 46 from a Complete NIRCam Grism Redshift Survey. 997 (2), pp. 207. External Links: Document, 2504.08028, ADS entry Cited by: §4.1.
- JADES: Differing assembly histories of galaxies: Observational evidence for bursty star formation histories and (mini-)quenching in the first billion years of the Universe. A&A 697, pp. A88. External Links: Document, 2306.02470, ADS entry Cited by: §4.1.
- The Luminosity Function and Star Formation Rate between Redshifts of 0.07 and 1.47 for Narrowband Emitters in the Subaru Deep Field. ApJ 657 (2), pp. 738–759. External Links: Document, astro-ph/0610846, ADS entry Cited by: §4.1.
- The evolution of dust attenuation in z 2-12 galaxies observed by JWST. Nature Astronomy 9, pp. 458–468. External Links: Document, 2402.05996, ADS entry Cited by: §2.
- New insight on the nature of cosmic reionizers from the CEERS survey. 685, pp. A3. External Links: Document, 2309.02219, ADS entry Cited by: §3.1.
- Deciphering stellar metallicities in the early Universe: case study of a young galaxy at z = 4.77 in the MUSE eXtremely Deep Field. 660, pp. A10. External Links: Document, 2111.14855, ADS entry Cited by: §3.1.
- The origin of scatter in the star formation rate-stellar mass relation. 484 (1), pp. 915–932. External Links: Document, 1805.05956, ADS entry Cited by: §1.
- JWST FRESCO: a comprehensive census of H + [O III] emitters at 6.8 ¡ z ¡ 9.0 in the GOODS fields. MNRAS 535 (1), pp. 1067–1094. External Links: Document, 2405.05111, ADS entry Cited by: §2.
- RASCAS: RAdiation SCattering in Astrophysical Simulations. A&A 635, pp. A154. External Links: Document, 2001.11252, ADS entry Cited by: §2.
- The hα and [o iii] λ5007 luminosity functions of 1.2 lt; z lt; 1.9 emission-line galaxies from hubble space telescope (hst) grism spectroscopy. The Astrophysical JournalA&AAJApJApJMNRASApJA&AApJApJApJApJMNRASApJApJMNRASarXiv e-printsA&AMNRASMNRASA&AMNRASApJMNRASA&AMNRASPASPA&AarXiv e-printsA&AMNRASApJA&AApJSApJApJMNRASMNRASMNRASApJMNRASMNRASMNRASMNRASApJMNRASApJApJARA&AApJARA&AApJMNRASThe Open Journal of AstrophysicsComputing in Science and EngineeringComputing in Science & EngineeringNatureApJJournal of Machine Learning Research 943 (1), pp. 5. External Links: Document, Link Cited by: §4.1.
- All the Little Things in Abell 2744: 1000 Gravitationally Lensed Dwarf Galaxies at from JWST NIRCam Grism Spectroscopy. pp. arXiv:2410.01874. External Links: Document, 2410.01874, ADS entry Cited by: §2, Figure 4, §4.1.
- Star Formation in AEGIS Field Galaxies since z=1.1: The Dominance of Gradually Declining Star Formation, and the Main Sequence of Star-forming Galaxies. 660 (1), pp. L43–L46. External Links: Document, astro-ph/0701924, ADS entry Cited by: §4.1.
- Scikit-learn: machine learning in Python. 12, pp. 2825–2830. Cited by: §5.
- IPython: A System for Interactive Scientific Computing. 9 (3), pp. 21–29. External Links: Document, ADS entry Cited by: §5.
- The JWST Emission Line Survey (JELS): an untargeted search for H emission line galaxies at z ¿ 6 and their physical properties. 541 (2), pp. 1348–1376. External Links: Document, 2410.11808, ADS entry Cited by: §1.
- RAMSES-RT: radiation hydrodynamics in the cosmological context. MNRAS 436 (3), pp. 2188–2231. External Links: Document, 1304.7126, ADS entry Cited by: §2.
- LyC escape from SPHINX galaxies in the Epoch of Reionization. MNRAS 515 (2), pp. 2386–2414. External Links: Document, 2207.03232, ADS entry Cited by: §2.
- The SPHINX cosmological simulations of the first billion years: the impact of binary stars on reionization. 479 (1), pp. 994–1016. External Links: Document, 1801.07259, ADS entry Cited by: §1, §2, §2, §4.3.
- The Luminosity Function and Stellar Evolution.. 121, pp. 161. External Links: Document, ADS entry Cited by: §3.1.
- The JWST EXCELS survey: probing strong-line diagnostics and the chemical evolution of galaxies over cosmic time using Te-metallicities. 540 (2), pp. 1800–1826. External Links: Document, 2502.10499, ADS entry Cited by: §1.
- Global Star Formation Rate Density over 0.7 ¡ z ¡ 1.9. ApJ 696 (1), pp. 785–796. External Links: Document, 0902.0736, ADS entry Cited by: §4.1.
- The Diversity and Evolution of Dust Attenuation Curves from Redshift z ~1 to 9. arXiv e-prints, pp. arXiv:2509.01795. External Links: Document, 2509.01795, ADS entry Cited by: §1, §2.
- A large H survey at z = 2.23, 1.47, 0.84 and 0.40: the 11 Gyr evolution of star-forming galaxies from HiZELS★. 428 (2), pp. 1128–1146. External Links: Document, 1202.3436, ADS entry Cited by: §4.1.
- The star formation main sequence and stellar mass assembly of galaxies in the Illustris simulation. 447 (4), pp. 3548–3563. External Links: Document, 1409.0009, ADS entry Cited by: §1.
- Re-evaluating old stellar populations. 479 (1), pp. 75–93. External Links: Document, 1805.08784, ADS entry Cited by: §2.
- Reconciling the Stellar and Nebular Spectra of High-redshift Galaxies. 826 (2), pp. 159. External Links: Document, 1605.07186, ADS entry Cited by: §3.1, §4.3.
- A SPectroscopic Survey of Biased Halos in the Reionization Era (ASPIRE): Spectroscopically Complete Census of Obscured Cosmic Star Formation Rate Density at z = 4–6. 980 (1), pp. 12. External Links: Document, 2412.06894, ADS entry Cited by: §1.
- Stochastic modelling of star-formation histories II: star-formation variability from molecular clouds and gas inflow. 497 (1), pp. 698–725. External Links: Document, 2006.09382, ADS entry Cited by: §1.
- Cosmological hydrodynamics with adaptive mesh refinement. A new high resolution code called RAMSES. A&A 385, pp. 337–364. External Links: Document, astro-ph/0111367, ADS entry Cited by: §2.
- Dust Attenuation, Star Formation, and Metallicity in z 2-3 Galaxies from KBSS-MOSFIRE. 871 (1), pp. 128. External Links: Document, 1805.00016, ADS entry Cited by: §1, §3.1.
- Decoding the variability in the star formation histories of z 0.8 galaxies. 539 (4), pp. 2891–2909. External Links: Document, 2504.05281, ADS entry Cited by: §1.
- The Variability of the Star Formation Rate in Galaxies. I. Star Formation Histories Traced by EW(H) and EW(HA). 892 (2), pp. 87. External Links: Document, 1912.06523, ADS entry Cited by: §1.
- Galaxy build-up in the first 1.5 Gyr of cosmic history: insights from the stellar mass function at z 4-9 from JWST NIRCam observations. MNRAS 533 (2), pp. 1808–1838. External Links: Document, 2403.08872, ADS entry Cited by: §2.
- Recalibrating the cosmic star formation history. MNRAS 490 (4), pp. 5359–5365. External Links: Document, 1910.05220, ADS entry Cited by: §1.
- The Evolution of the IR Luminosity Function and Dust-obscured Star Formation over the Past 13 Billion Years. 909 (2), pp. 165. External Links: Document, 2101.04734, ADS entry Cited by: §1.